OpenStudy (anonymous):

Click here if you want to help me with Alg 2 study guide for the test. I'll take any help. My math teacher sucks

OpenStudy (anonymous):

Fire away, I'm only algebra 1 but it's an advanced class so maybe I can help

OpenStudy (anonymous):

I looked at the first one...

OpenStudy (anonymous):

Juanita's here we are saved LMFAO

OpenStudy (anonymous):

she gave you an equation and a point of (4/3, 0)

OpenStudy (anonymous):

why 0?

OpenStudy (anonymous):

so basically set the equation to 0 and solve for x (we know the answere will be 4/3 but we must prove

OpenStudy (anonymous):

oh I was looking at the wrong problem

OpenStudy (anonymous):

Just give her the dang medal already lol

OpenStudy (anonymous):

4/3 is x value and 0 is y value.. so normally you are use to seeing y = x + 2 will if I told you y = 0 and solve for x

OpenStudy (anonymous):

0 = x - 2 2 = x

OpenStudy (anonymous):

perfect - now do you understand why she set it to 0? she move all terms to the left side

OpenStudy (anonymous):

I think theyre already on the left side and equal to 0

OpenStudy (anonymous):